OpenStudy (anonymous):

If f(x) = x - 1 and g(x) = 2x - 3, find (fg)(x). Path: pWords:0 3 points QUESTION 7 7. If f(x) = x - 1 and g(x) = 2x - 3, find (g - f)(x). Path: pWords:0 3 points QUESTION 8 8. If f(x) = x + 4 and g(x) = 5x - 2, find (g-f)(2). Path: pWords:0 3 points QUESTION 9 9. If f(x) = x + 4 and g(x) = 5x - 2, find (g + f)(-3). Path: pWords:0 3 points QUESTION 10 10. If f(x) = x + 4 and g(x) = 5x - 2, find (gf)(4). Path: pWords:0 3 points Click Sav

OpenStudy (astrophysics):

Hey try to limit 1 question per post, lets look at #6/ first one you have up there. \[(fg)(x) = f(x)g(x) = (x-1)(2x-3)\] this means you multiply the to the functions

OpenStudy (astrophysics):

two*

OpenStudy (astrophysics):

\[(g-f)(x) = g(x)-f(x) = (2x-3)-(x-1)\] then your next one means subtract the two functions

OpenStudy (astrophysics):

I think you should be able to do the rest, don't forget to simplify!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

are they all the same setup?

OpenStudy (astrophysics):

Yeah pretty much, just the sign might be different

OpenStudy (astrophysics):

Like (g+f)(x)

OpenStudy (astrophysics):

That just means g(x)+f(x)

OpenStudy (anonymous):

ok I'm still having trouble finding the answer

OpenStudy (astrophysics):

For the first one you're just distributing you know how to do that?  the arrows indicate multiply